OpenStudy (anonymous):

Given the wavelength of the corresponding emission line, calculate the equivalent radiated energy from n = 3 to n = 2 in both joules and electron volts. Also calculate the frequency of the wave.

OpenStudy (anonymous):

energy in at the nth level = - A/n^2 look up A [13.something eV?] make transition from n=3 to n=2 and take that energy = h c / lambda, h is Planck's, c is light velocity, lambda is wavelength

OpenStudy (anonymous):

Just one question, what does A stand for?

OpenStudy (anonymous):

A is a constant and is equal to -13.6eV. It's included in the following equation:\[\frac{ hc }{ \lambda }=-13.6\left( \frac{ 1 }{ n _{1}^{2} }-\frac{ 1 }{n _{2}^{2} } \right)eV\]where n1 is the initial electron energy level and n2 is the final electron energy level. That constant is known as the Rydberg constant and is also the ground state energy for the hydrogen electron.